40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog, Page 7 of 60  Not logged in ELOG logo
ID Date Author Type Category Subject
  2731   Fri May 7 13:15:07 2021 shrutiUpdatePSOMACavity locking, feedback to laser

[Shruti, Rana]

I made that the ITC 502 current driver was in constant current mode where the feedback coeff is 20 mA / V \pm 5% and bandwidth 500 kHz; with the laser coefficient being 100 MHz the net feedback coefficient would be 2 GHz/V.Then, I looked at the level of the PDH signal which was roughly \pm 1V.

In order to attenuate the signal going in to the MOD IN of the ITC 502 current driver, Rana suggested using the 50 Ohm output of the SR560 with two series 20 dB attenuators and a BNC T with one port going to the MOD IN and the other port terminated with 50 Ohm.

With the SR 560 set with 1 Hz low pass and gain of 10, the cavity locked for 3-10 seconds when it came close to resonance as seen on the monitor.

The three attachments are the PDH signal (purple), cavity reflection (blue), and cavity transmission (yellow) when the cavity seemed to lock.

  2730   Thu May 6 10:48:40 2021 aaronDailyProgressLaserdelay line frequency discriminator

delay line measurements

All of the measurements above had moku presenting a 50 Ohm load. Maybe I was too close to the 1 dB compression point, typical output powers were 0-20 dBm and the 1 dB compression point is 20-30 dBm for most of the above.

I'll repeat the steps here with some modification, since the original didn't make it into the text

  1. Check the RF power of both legs before the mixer, and ensure the LO is at +7dBm and the RF is within the linear range of the mixer.
  2. Calibrate the frequency discriminator by tuning the beat note to the upper and lower null fringe. This tells us the amplitude of the sine we are measuring, which will be the same as the slope of the sine at the half fringe.
  3. Tune the beat note to half fringe by changing the laser current until the frequency is halfway between the upper and lower null frequencies. I think this is preferable to choosing the DC voltage to be halfway between the voltage nulls, since there's some current-to-power coupling.
  4. Record several spectra using the SR785 and SRmeasureWideSP.py
laser pair lower null frequency lower null DC voltage upper null frequency upper null DC voltage voltage at half fringe frequency at half fringe RMS voltage at half fringe Hz/V calibration start time of measurement drift during measurement notes
Rio N x Rio S laser 63.3 MHz 0.15 V 100.6 MHz -3.64 V -1.27 81.9 MHz 70 mV 3.12 MHz/V Thu May 6 12:03:23 2021 7 MHz  
Rio N x Teraxion laser 64 MHz -0.3 V 98 MHz -3.1 V -1.7 V 84 30 mV 4.55 MHz/V Thu May 6 17:37:31 2021   added 3dB attenuator to LO path, switched to ZHL-1A
Rio S x Teraxion laser 68 MHz -0.5 V 100 MHz -3.2 V -1.6 V 84 7 mV 7.55 MHz/V Thu May 6 19:01:25 2021   3+6 dB attenuators before LO

 

It turned out I simply needed to reset the gpib connection for the script to run. Data collection running smoothly now. The calibration is

\frac{\mathrm{Hz}}{\mathrm{V}} = \frac{\Delta f}{\Delta V} * \frac{2}{\pi}

 


Spent some time tuning the N and S lasers to what seemed like good temp/current regimes... but after doing so for the S x Teraxion beat note, I'm seeing a shockingly small rms voltage at the half fringe! In particular, some distinctly periodic noise disappeared appears in both N and S laser beat notes, but is smaller for the S laser. Attachment 1 is that noise for the S laser. Some possibilities

  • the ITC510 is just noisier than the ITC502. This does seem consistent with the current noise measurement we took last week, and the periodic noise on the beat note looks similar to that on the current. [add evidence please]
  • I haven't actually put the N laser into a low noise operating point. Need to systematize my process for setting T and I.
  • The N laser is just noisy. Obviously the least desirable of these.

Another aside, I do think I've been approaching the saturation limit of the 1611, and that's the most likely culprit behind the harmonics I've been seeing. The SX beat note measurement was around 700 uW, and the PD saturates at 1 mW.

 


update: the results from these delay line measurements are attached. Obviously still some issues, including regions of negative estimated laser noise where we need more averaging, and I don't really believe either of the current noise measurements at this point.

  2729   Wed May 5 22:21:17 2021 ranaDailyProgressGeneraldelay line frequency discriminator

no need to eval them all, just use the ZHL-3A. Make sure you're not saturating by checking the datasheet. ask me if you're not sure. For example these things can get damaged if the Moku presents a non 50 Ohm load. You have to make sure the output level is well below thi 3rd order intercept point.

 

  2728   Wed May 5 11:54:19 2021 aaronDailyProgressLaserdelay line frequency discriminator

I measured the NX beat last night with the delay line frequency discriminator and an Agilent 8447A amplifier. However, this morning when I looked at the beat note spectrum, I found that the amplifier was generating many harmonics of the resonance. While looking for a working amplifier, I found more oscillations, so decided to test them systematically...

testing RF amplifiers

When powering up these amplifiers, I'm carefully following the steps below. When powering down the amplifiers, I reverse the steps. I'm supplying floating DC voltage with a GW Instek GPS-3030D power supply.

  1. connecting output to 50 Ohm load
  2. Supplying DC power
  3. Connecting RF signal to input

Here are my findings

amplifier DC supply voltage gain noise figure (datasheet)
VSWR
in // out
oscillating? strongest harmonic frequency attachment notes
ZX60-100VH+ +12 V 36 dB 4 1.6 // 1.5 yes 2 * f_beat 3  
ZHL-3A +24 V 24 dB 5.2 2 // 2 yes

2 * f_beat

1 * f_beat only appears above 8 MHz

1  
ZHL-1A +24 V 16 dB 8 1.28 // 1.27 no   2  
ZHL-2010+ +12 V 20 dB 3.98 2 // 2 yes, somewhat 2 * f_beat 4 I should have been driving this amp above 50 MHz
LNBA-0-13-AB-68.85 +15 V 13 dB -170 dBc/Hz input referred noise @ 10 kHz < 2 yes 2 * f_beat 5  
  • The gain, noise figure, and VSWR are all from the datasheet, not from my measurements. In all of the cases where the amplifier generated some harmonics, there were no spuri ous peaks without some RF signal connected to the amplifier input.
  • I noticed that on some amplifiers, when the beat note was below 8 MHz, the harmonic was at 2 * f_beat, while above that frequency a second harmonic appears at 1 * f_beat. Could have just been below the noise floor. In all cases, the second harmonic was higher than the first harmonic
  • I tried switching to a shorter cable between the amplifier output and the Moku spectrum analyzer and then between the 1611 and amplifier input, but didn't see any difference.
  • We also have a second, BNC-terminated ZHL-1A, but one of its DC supply filter capacitors is not fully soldered so I didn't test it.

This behavior is so consistent, I think it must be expected in a functioning RF amplifier, a simple error in my measurement setup, or a very common failure mode of the amplifiers.

Delay line frequency discriminator

Despite the minor mystery above, at least some of these amplifiers should be suitable for the delay line frequency discriminator measurement. Attachment 6 is my measurement setup today.

 

 

Aaand the elog glitched out and deleted my entry. Fortunately got screen shots.

  2727   Wed May 5 00:27:10 2021 ranaDailyProgressLasermarconi locking

Huzzah!

The spectrum looks curiously high though. I think the noise should be close to 1e4 * (1/f) Hz/rHz. May just be you hove to set the 785 to AC coupled and then have it auto-range the input ~20-30s after the loop is locked. The AC coupling is at 0.15 Hz, so it needs a few e-foldings to settle.

But the low frequency noise looks right, so the setup seems to be working correctly.

 

  2726   Tue May 4 10:50:22 2021 aaronDailyProgressLasermarconi locking

I was able to acquire lock after swapping the LB1005 server controller for the SR560 ((10 MHz BW compared to 1 MHz), and an SLP-1.9+ instead of the 10 MHz LP. On the LB box I played around with the PID parameters and eventually got these to work best

  • LF sent to channel A, channel B is terminated with 50 Ohms.
  • PI corner at 10 Hz
  • Gain of 10, LF gain limit of 30 dB
  • I set the input offset by turning the LB box to LFGL mode and tweaking the offset until the spectrum analyzer showed the LO was stably at the beat note frequency
    • Confirmed that this offset also sends the average error output to 0 V with the LO turned off, so it's a sensible offset

I sent split the modulation input (400 kHz/V) of the Marconi with a T junction, and am measuring the spectra with SR785 and Anchal's handy wide spectrum script in labutils/netgpibdata. I'm monitoring the RF and LO notes on moku to make sure the lock is maintained throughout the measurement.

The first spectra I took of the NS beat note (fig 1) showed a high noise floor starting at 10 Hz, probably because I've got no extra 'I' gain above that frequency. And, after switching to the NX beat, I can't acquire lock (Teraxion over fiber seems a bit noisier than the Rio S laser).

I'm logging 30 min of data measuring the NX beat with the delay line frequency discriminator (10x gain on SR560 amplifier, with 3 kHz LPF). Will finish up the other two beat notes in the morning.

 

  2725   Mon May 3 23:28:56 2021 ranaDailyProgressLaserdelay line frequency discriminator

When the beat moves a lot, it may be easier to use a passive device rather than a PLL.

This gadget just requires a few RF parts and a long cable. Takes less than an hour to get it working.

  2724   Fri Apr 30 15:15:33 2021 ranaDailyProgressLasermarconi locking

I think that mixer's IF output is DC coupled, so you should not use an offset. Ideally, when you bring the LO close to the beat note freq, the PLL will catch the lock by feeding back to the Marconi's DC coupled FM Dev input, such that the LO and the RF have the same frequency, but are shifted by 90 deg.

I suggest using a 5 MHz LP, if you have one.

SR560 should be DC coupled - see Shruti's 4om elog on PLL locking.

  2723   Fri Apr 30 12:07:25 2021 aaronDailyProgressLasermarconi locking

[aaron, shruti]

I set up the Marconi to measure the frequency noise on the NS beat note, with both lasers controlled by the commercial Thorlabs drivers. Shruti helped me get everything locking... relative to my setup with a function generator

  • I needed an SMA-T terminated with 50 Ohms after the mixer to handle the impedance matching at high frequency
  • required a slight offset of the carrier relative to the beat note to keep the IF above the mixer's low frequency cutoff.
  • The LF from the Marconi is picked off of the internal modulator before adding the external modulation, so the error signal should be read off instead from a low-passed version of the IF. We picked off the error signal with a BNC T junction before the high-passing SR560, so we preserve the DC error signal but don't inject it back into the carrier.
    • This still isn't quite making sense to me. I'll attach some sideband pictures to see if I've got it right.
  2722   Thu Apr 29 17:26:24 2021 aaronThings to BuyGeneralfiber box

fiber box

I looked into boxes for fiber components so we can eventually add some acoustic damping around our lasers. There are some pretty good rack-mounted breadboard options on Thorlabs, which I've added to the PSOMA hardware inventory.

What is a good type of foam that won't degrade over time? Thinking we want something like this pick-and-pack so we can modify the shape to fit our application after fixing the fiber components.

This does raise the question of how much phase noise to expect in transferring the lasers by fiber from the rack to a fiber launch on the table, and whether the vibrations of the rack and its electronics would make some noise. I expect we'll actually want to lock the fibers to the cavity independently... does this mean we don't actually need a fiber PD to mix the lasers before they launch into the interferometer?

RF electronics

I've added list of RF electronics in cryo lab (but not the ones we're still waiting on from the order) to the PSOMA hardware inventory. All quantities are lower bounds, since I only counted what was in the stock drawer.

  2721   Thu Apr 29 17:25:55 2021 aaronNoise HuntingLaserphasemeter drift

Continuing this, I measured the spectrum of the NS beat note over 20 min (10 min x 2) with persistance on to see how much the note drifts. I observed about 1 MHz/min drift on moku spectrum analyzer (attachment 1).

Afterwards, I measured the same beat note on the phasemeter over 5 minutes. Indeed, it appears the phasemeter loses lock after several minutes (attachment 2), whereupon it drifts by again 1 MHz/min. (attachment 2). I did notice this behavior on some of the data I took last week, but couldn't explain it and attriubuted it to the phasemeter settling in (which in retrospect doesn't make sense, because the BW is kHz while the 'settling' would have been for tens of seconds).

 

Rana and I measured the current noise of the ITC510 using SR560 + moku, both driving the laser diode and a 20 Ohm dummy load. We noticed some unusual-looking noise on the moku display, and I'll post some details when I've had a chance to plot the data.

Update:

Attachment 1 shows the measurement setup. In the upper diagram, we've connected the ITC502 to the N RIo laser using the usual DB9 cable, but with a breakout board at the driver and clips sent to a high impedance voltmeter on moku.  In the lower diagram, we've replaced the DB9 cable with two wirebound resistors stuck in the DB9 connector of the breakout board (50 Ohm across pins 1,5 for the interlock; 20 Ohm across pins 3 and 7 to simulate the diode).

Update:

The current noise as measured by the Moku spectrum analyzer while driving the N laser is in attachment 2. Obviously it's high... I'll start adding this curve to future beat note measurements involving the ITC510. To get the current amplitude spectral density from the Moku's reported power in dBV, I divided by the sensing resistance (which I took to be 20 Ohm) and resolution bandwidth of the spectrum analyzer. I'm not sure this is the correct sensing resistance -- the laser datasheet shows 20 Ohm parallel, not in series, with the current drive. x-axis in Hz.

I_\mathrm{ASD} = \frac{10^{V_\mathrm{dBV}/20}}{R_\mathrm{sensing}*\sqrt{RBW}}

  2720   Thu Apr 29 13:01:13 2021 ranaComputingGeneralcaps lock -> ctrl

For conda, its possible to export a yml file with the requirements based on your current env. That should allow you te setup identical envs. YMMV.

For tmux, similarly, you can change conf settings in tmux.conf. And save window/session stuff.

 

  2719   Wed Apr 28 15:26:10 2021 ranaNoise HuntingLaserphasemeter drift

my guess exactly. I will ask Moku people about it, but my guess is that we need to do a traditional phase lock using some mixers, etc like what Shruti had working at the 40m. I suggest scrounging some parts from somewhere in WB / 40 for this week, but fill up our purchase spreadsheet with some RF shopping list that we ran go over in our meeting tomorrow (Thurs)

Quote:
 

I didn't replicate 'MHz drift over several minutes' exactly, but I suspect our beat note is pushing the BW limit of the phasemeter.

nice idea with the triangle wave. I was thinking of putting some white noise into the FM dev input of the Marconi, but your way is easier. Will definitely need Marconi for the PLL setup.

  2718   Wed Apr 28 12:09:28 2021 aaronNoise HuntingLaserphasemeter drift

Diagnosing the phasemeter

I took these diagnostic steps:

  1. Drive moku spectrum analyzer and phasemeter with a bandpassed (SBP-30+) -30 dBm sine at 29 MHz from a DS345 function generator
    • Spectrum analyzer shows rock solid beat note, as expected (attachment 1)
    • The moku phasemeter frequency estimate shows no drift at the level of the ~3-4 Hz rms noise. Phasemeter settings for attachment 2:
      • Auto frequency
      • 10 kHz BW
      • AC coupled, 50 Ohm input impedance
      • 10 Vpp range
      • 488 Hz acquisition speed
    • With 1 Vpp range selected, still no drift above the noise
    • With 1 Vpp and a 10 Hz bandwidth, there is at most a few 100 mHz drift over 4 minutes (attachment 3)
    • With 1 Vpp and 600 mHz bandwidth, the phasemeter at first has trouble acquiring lock, but then shows the same at most couple 100 mHz drift over a couple minutes (attachment 4)
  2. Drive moku phasemeter with an unfiltered -10 dBm sine at 29 MHz from a DS345
    • The spectrum is full of features. 29 MHz is close to the limit of the DS345. (attachment 5)
    • The phasemter settings for attachment 6 show no drift over several minutes
      • Auto frequency, AC coupled, 50 Ohm input impedance, 1 Vpp range, 488 Hz acquisition speed
      • 10 kHz BW
  3. Drive moku spectrum analyzer and phasemeter with a bandpassed (SBP-30+) -25 dBm sine at 29 MHz, with a 143 Hz triangle wave FM of 1 MHz deviation
    • Spectrum is attachment 6
    • At 10 kHz bandwidth, the phasemeter has no trouble tracking the true drive frequency (attachment 7) and does not exhibit drift above the 1 MHz modulation.
      • Settings other than bandwidth are still Auto frequency, AC coupled, 50 Ohm input impedance, 1 Vpp range, 488 Hz acquisition speed
    • At 600 Hz BW, the phasemeter estimates the frequency a bit low (28.8 MHz) and with only about 40 Hz deviations
    • At 10 Hz BW, the behavior is inconsistent. Sometimes, the phasemeter exhibits a 20 Hz/min drift (attachment 8). Reacquiring lock sometimes results instead in some few Hz noise around a fixed frequency (attachment 9).
  4. Drive moku with a bandpassed (SBP-30+), -25 dBm sine at 27.5 MHz, with a 143 Hz triangle wave FM of 2 MHz deviation
    1. At 10 kHz BW (other settings same as before), the phasemeter initially drifts rapidly (MHz over seconds), then stops below the center frequency (attachment 10)
    2. At 40 Hz BW, the same behavior is visible over a somewhat smaller frequency range and longer timescale (10s Hz drift over 10s seconds, before stopping at a fixed frequency) (attachment 11)
    3. At 600 Hz BW, there is not drift and the frequency is steady at but underestimated by 0.5-1 MHz.

 

I didn't replicate 'MHz drift over several minutes' exactly, but I suspect our beat note is pushing the BW limit of the phasemeter.

  2717   Wed Apr 28 10:29:29 2021 aaronComputingGeneralcaps lock -> ctrl

How about caps lock to escape? That's my usual.

On spirou or gaston, there are a few virtual envs I typically use...

  • liso -- for liso
  • moku -- for interfacing with moku
  • psoma -- for using scripts in the PSOMA repo, including qlance and some plotting tools

I usually also have a couple tmux sessions running for quick access and segmentation, such as

  • ipython -- for the kernel, which I usually launch from a psoma env
  • moku -- has a moku env

Any tips on good ways to quickly / automatically instantiate or document some of these?

  2716   Wed Apr 28 09:46:12 2021 ranaNoise HuntingLaserproblems with the phasemeter

The Moku phasemeter seems to be slipping its phase.

While looking at the SX or NX beats, we noticed that there was a steep linear trend in the tracked frequency. About 2 MHz/min.

However, just plotting the spectrum and watching the peak frequency v time (Shruti's elog attachment), we see that the frequency drift is actually not real.

So, what's going on?

  2715   Tue Apr 27 20:24:18 2021 ranaUpdatePSOMAPDH error signal

aww, such a milestone deserves a longer movie! At least a few more flashes. And maybe a soundtrack?

Quote:

Attachment 1: PDH error signal with monitor flashes

  1. The PDH signals are surprisingly large, but perhaps it makes sense if we calculate the optical and electronic gain.
  2. Should adjust the LO phase (by addinng / subtracting cable length) to maximize the peak-peak signal (after confirming that the 1811 is not saturating).
  3. Figure out where we should actuate to lock the laser to the cavity.
  4. Is it time to revive a custom current driver?
  2714   Tue Apr 27 20:19:09 2021 ranaUpdateLaserlaptop broken

The windows laptop that we were using to control the Teraxion laser has got a corrupt disk and the Winows autofix feature is failing to fix it after a couple hours. So we need to get another Windows laptop and install the software on it.

My guess is that the power was pulled on this without clicking shutdown. Since this computer has no battery, the sudden power loss may have caused the disk corruption.

  2713   Tue Apr 27 19:37:49 2021 shrutiUpdatePSOMAPDH error signal

Attachment 1: PDH error signal with monitor flashes

Using the setup shown in Attachment 3 (diagram here) (cavity design here), where the RF output of the free-space 1811 photodiode is sent to a Level 17 mixer ZFM-3H-S+ and then to a low pass filter SLP-1.9+ (cut-off at 1.9 MHz), I observed the PDH error signal on an oscilloscope while scanning the temperature.

Attachment 4 is a single-shot measurement of the error signal.

EOM: MPX-LN-0.1 driven with -5.8 dBm at 33.59 MHz using the OCXO pre-amp EOM output at 30 dBm attenuation.

LO: Driven at 33.59 MHz using the LO output of the OCXO pre-amp with no attenuation, estimated to be ~13.8 dBm.

The power levels of the OXCO outputs were the peak powers calculated by the Moku while observing the spectrum. The Moku itself produces a lot of distortion (many prominent harmonics were seen) while measuring this spectrum, these harmonics were less prominent when the scale was changed or when the input was attenuated sufficiently.

Attachment 2:

Picture of beat note measurement between North and Teraxion requested by Rana.

  2712   Tue Apr 27 18:08:05 2021 ranaComputingGeneralcaps lock -> ctrl

remapped caps lock to CTRL on Spirou, because its nicer that way for emacs, and no one ever needs CAPS LOCK for anything (unless you are on twitter)

ALSO, I recommend not using the base env in anaconda, but rather make a dedicated env called `lab` or something. Leave base for an unmodified python instance. all the fancy NDS stuff in `lab`

  2710   Tue Apr 27 14:15:40 2021 aaronUpdateLasermeasuring NX beat note

Turned on the Teraxion laser, and am repeating the beat note measurements.

Also picked up some minicircuits components and polaris BS mounts from Downs.

  2709   Mon Apr 26 20:13:24 2021 aaronUpdateLasermeasuring beat notes

Not great data from these. Attachment 1 contains the several beat note frequency noise ASD I measured, mostly last week. You can see that the noise on the SX beat is two orders of magnitude higher than the others, which doesn't make sense because the N laser driver is noisier than the S laser driver. 

Attachment 2 shows the resulting three corner hat, where I've simply averaged the median estimate for each beat note separately then applied the rotation to get the individual lasers. The estimates for S and Teraxion lasers are identical and the estimate for N is always negative, because the sum is always dominated by the SX beat measurements. Obviously not to be used or trusted. 

Perhaps I'm just measuring the noise of the moku phasemeter in 'unlocked' state, since the beat notes are all well below my expectation of ~1e4 Hz/rtHz @ 100 Hz. Unfortunately will need to take new measurements, or take a closer look at the raw timestreams. 

Note on the data and script:

  • All raw data are available on gaston and spirou in controls:~/cryo_lab/data/3CH/*note*.npy
  • The beat note spectra over 7 minutes of data are estimated with Welch's method performed with no overlap and median averaging. The data are transformed 'decade by decade' using as many averages as possible in each frequency decade, as in labutils/moku/modifiedPSD.
  • The plots were generated from ThreeCornerHat.ipynb, available on git.ligo in cryo_lab/scripts. 
  • The data for the beat note spectra plotted below are on git.ligo (lfs) in cryo_lab/data/3CH/notes.npy. There is also an array beats.npy in the same folder, containing the key for which beat note corresponds to which pair (0->NS, 1->NX, 2->SX)
  2708   Thu Apr 22 20:25:49 2021 shrutiUpdatePSOMAsetup to do both locking and incoherent 3CH

Some updates to the setup...

Quote:

Highlighted part of the setup shows what is required to continue with the cavity work. The 10% beam from the south laser and other available items should be sufficient for Aaron to continue with the incoherent 3CH stuff.

 

  2707   Thu Apr 22 12:56:18 2021 ranaUpdatePSOMACavity aligned and resonant flashes observed

smileyheartlaugh

Quote:

After aligning the cavity roughly to the solution and adding lenses according to the mode-matching solution found in elog 2703, I observed some higher order mode flashes on the camera and also saw dips in the reflected power (as in the first few seconds of Attachment 1). I adjusted the cavity yaw alignment since the higher orders were along the yaw-axis while Aaron tuned the temperature sweep and offset to see TEM00 flashes (as in the last few seconds of Attachment 1).

Mode matching efficiency ~ 73%, since the dips corresponded to 150 mV and the non-resonant region to 550 mV as seen on the oscilloscope.

 

  2706   Thu Apr 22 12:51:15 2021 aaronUpdateLasermeasuring beat notes

Measuring the SX beat note again while the cavity is not flashing. I tuned the current of the S laser from __6.63 mA

  2705   Thu Apr 22 10:43:25 2021 shrutiUpdatePSOMACavity aligned and resonant flashes observed

After aligning the cavity roughly to the solution and adding lenses according to the mode-matching solution found in elog 2703, I observed some higher order mode flashes on the camera and also saw dips in the reflected power (as in the first few seconds of Attachment 1). I adjusted the cavity yaw alignment since the higher orders were along the yaw-axis while Aaron tuned the temperature sweep and offset to see TEM00 flashes (as in the last few seconds of Attachment 1).

Mode matching efficiency ~ 73%, since the dips corresponded to 150 mV and the non-resonant region to 550 mV as seen on the oscilloscope.

  2704   Wed Apr 21 15:26:36 2021 aaronUpdateLasermeasuring beat notes

I let the NS beat run this morning to afternoon. Switched to the SX beat around Wed Apr 21 15:33:37 2021. This required me tuning the S laser current, so please note the time in the elog (or better yet stop the measurement script on spirou tmux session) if you change the laser current for further locking activities.

  2703   Wed Apr 21 09:52:11 2021 shrutiUpdatePSOMACavity design

Latest cavity design (attachment 2) and mode-matching solution (attachment 1).

Attachment 3 suggests the tolerance of this setup and attachment 4 is an alternative mode-matching solution.

  2702   Tue Apr 20 19:09:37 2021 aaronUpdateLasermeasuring NX beat note

Set up a script (cryo_lab/scripts/phaseMeter_loop.py) to repeatedly measure the North Rio laser x Teraxion laser beat note, and left running in a tmux session on spirou.

  2701   Thu Apr 15 19:24:35 2021 shrutiDailyProgressPSOMARoC measurement

[Rana, Shruti]

Attachment 1: Figure updated with beam profiles measured for the concave side of the optic. Consistent with RoC = 1 m concave.

  • Turns out the optic was installed on the mount the with the side with HR coating for 1550 nm at the back. Verified by observing the color of reflected light from both sides: the concave side reflected green light while the convex side reflected red light. The side HR coated for 1550 nm must be AR coated for 1550/2 nm (close to visible red);  the concave side showed no red reflection and therefore is consistent with being HR coated for 1550 nm.
  • The RoC measured for the convex side is not -1m because of the presence of additional material with a different refractive index in the path.
  • The blue circles in the plot with x-values of 1 m and 0.85 m are not accurate because they were measured at large incidence angles on the curved optic. The remaining values were measured at small incidence angles. The results therefore seem consistent with a radius of curvature of 1m.

Attachment 2: MATLAB code to generate plot

Attachment 3: Beam profile data

 

This refers to curved optic labeled SN 1.0-Si-1.0M; previous record of it found in elog 2562.

  2700   Thu Apr 15 16:28:20 2021 shrutiUpdatePSOMAsetup to do both locking and incoherent 3CH

Highlighted part of the setup shows what is required to continue with the cavity work. The 10% beam from the south laser and other available items should be sufficient for Aaron to continue with the incoherent 3CH stuff.

  2699   Wed Apr 14 17:15:23 2021 shrutiDailyProgressPSOMARoC measurement

Attachment 1:

Result. Measured radius of curvature seems to be closest to 0.4 m convex.

 

Attachment 2:

Beam profile bitmap images. Widths taken read from the image.

 

Attachment 3:

Code used to plot Attachment 1.

  2696   Tue Apr 13 11:31:37 2021 aaronDailyProgressGeneralsome parts to W bridge

enter: Tue Apr 13 11:31:35 2021

Picked up some boxes from Downs:

  • pin connectors for our EOM
  • 0.5" optic mount from newport
  • some batteries and battery holders for EE

Mostly just math.

exit

  2695   Thu Apr 8 22:47:03 2021 shrutiDailyProgressPSOMACavity locking and temperature scan

[Rana, Shruti]

Today Rana pointed out several improvements to the setup including fixing some finicky knob situations. The cavity and setup are now in a good state with monitor and PD connected, although we still have not seen resonance (Will be Attachment 2 soon).

According to the calculations in Attachment 1 a 1V change in the 'TEMP TUNE' input of the ITC-502 temperature (and laser current) controller would correspond to a change of 1 GHz in laser frequency at the current operating point (8 kOhms 'T_ACT' as displayed).

We send in a 1 Hz triangle wave with V_pp=1V into the TEMP TUNE input and make tiny changes to the alignment while monitoring the monitor for flashes and the DC reflected power for a steep drop.

RXA edit:

  • We measured the input coupler transmission to be 2.5% for S-pol, so the cavity Finesse = 2*pi/(1 - T) ~ 250
  • From the overhead picture, you can see that the cavity length is ~15-16 screw holes, so L_{RT} = 16" = 40 cm. So the FSR = c/0.4 ~ 1 GHz.
  • So, assuming the cavity linewidth is then FSR/Finesse = 3 MHz, we want to make sure that our scan velocity is no faster than ~ 1 GHz/s, so that we are able to see the resonances easily on a 1 s oscope trace.
  • The diamter of the beam transmitted through the curved mirror of the cavity is ~3 cm on the monitor.
  • The mode-matching is still far off, since the beam after one round trip is much bigger than the input, but I estimate the mode matching is ~30-50%. Good enough for getting some flashes.
  • It would be good to find a mount for the 1" optics that allows us to cleanly get the 4-ports without clipping.

Other To Do

  • Get a new power adapter for the camera (weird flashes occur when shaking it, confuses the alignment/locking process)
  • Find/purchase a thread adapter to attach longpass filter to Watec camera
  • Eventually organize and safely store all ND filters and other optics in the cabinet
  2694   Tue Mar 30 15:00:56 2021 aaronHowToGeneralWhere is three corner hat data? Cryo repos?

Code and data for the cryo lab are in git.ligo.org within the Cryo Lab group

  • The data repo holds version controlled data, which should all be added using git lfs. File names should contain a date or other identifier that maps onto an associated elog entry.
  • The scripts repo contains scripts and notebooks for analyzing data
    • The notebook ThreeCornerHat.ipynb is what I used to generate three corner hat plots above
    • The notebook currentNoise.ipynb is what I used to look at the current noise.
    • Once I'm satisfied with these, I'll make a .py file that performs the measurement and generates plots in one go and add the useful sub-routines to labutils
  • Moku scripts
    • To take these measurements, I've been using the scripts in labutils/moku
    • The moku repo also includes liquid instruments' lireader repo for converting their proprietary .li binary format to a numpy object.
  2693   Tue Mar 23 11:07:56 2021 shrutiDailyProgressLab Worklab temperature logging

Lab temperature is now stabilized. The particle counter is now logging the temperature, but not the AD590 which records X1:AUX-LAB_TEMP_F.

The AD590 history also seems different from the previous record.

  2692   Mon Mar 22 21:41:09 2021 ChrisDailyProgressLab Worklab temperature logging

standalone_edc now works, and data from the temperature sensors (and other aux epics channels) is again being recorded.

The version of standalone_edc we had from RCG 4.0 silently fails to acquire data, unless symmetricom timing hardware is present. We don’t have such hardware in the cymac. I upgraded cymac1 to RCG 4.1 in order to get a newer version of standalone_edc. It can now use the IOP model as a timing source (--sync-to=x1iop_daq was added to the argument list). All models were rebuilt and reinstalled after the upgrade.

To minimize the chance for confusion, I removed the old nds/daqd service files, and archived the /opt/rtcds/tst/x1/target/fb directory. None of that is used anymore.

  2691   Mon Mar 22 13:28:32 2021 aaronDailyProgressLab Worklab temperature logging

Attachment 1 shows the most recent temperature data available on nds.

Separate from the dust monitor, we have an AD590 set up as a temperature sensor read by the Acromag, which interfaces with cominaux. This is the channel that cymac is supposed to write to frames as X1:AUX-LAB_TEMP_[K, F, C], but the value read by cominaux is currently (and has been) 0. I've confirmed the signal is reaching the acromax ADC on the correct channel.

The dust monitor is a particle counter (GT-526) that also shows temperature and relative humidity, and is read by cominaux over ethernet. Despite an "error reading and converting data from the counter" shown in cominaux:~/services/particlecounter.log between Feb 3 and today, caget'ing the channel C5:PEM-COUNT_TEMP (or the particle counts) returns the expected value (89-90 F). However, cymac has not been logging these frames and I can't access live or lookback data through nds.

I see that daqd.service and nds.service on cymac1 are both dead linked to /etc/systemd/system/[nds, daqd].service. Neither service script still exists in cymac1:/etc/systemd/system, only the 'old' version of these scripts. I think binaries for nds and daqd now live in /opt/rtcds/tst/x1/target/fb, and the standalone_edc service is meant to pull data from epics.

If I run standalone_edc, I get an error 'edc.ini' failed to open, and there are no channels to record; this service is not running on cymac1. /opt/rtcds/rtscore/edc_status.json is empty.

Quote:

What happened to the temperature logging? Has the CDS system stopped recording data due to model changes, or is it just the python for the dust monitor?

 

  2690   Mon Mar 22 12:05:41 2021 ranaDailyProgressLab Work 

What happened to the temperature logging? Has the CDS system stopped recording data due to model changes, or is it just the python for the dust monitor?

  2689   Sun Mar 21 12:52:43 2021 ranaHowToNoise Budgetphase noise

Youtube on phase noise in osc (https://youtu.be/wByzymJ0Ppc)

  2688   Fri Mar 19 11:08:15 2021 aaronDailyProgressLab Work 

Lab temperature is high again (85 at the particle counter), but apparently hasn't been logging. Attachment 1 is the set of status screens -- still no GPS, and yellow CFC.

incoherent three corner hat

Measured beat note with 3 lasers incident on 1611, RF out sent through SLP-21.4 to moku phasemeter

  • python [path_to_scripts]mokuPhaseMeterTimeSeries.py -c ch1 -d 450 -i 10.0.5.220 -s veryfast --useInternal --fileType bin --altFileName "L1L2_note"
    • where L1L2 are the lasers being measured (SX, NX, NS)

  • DC level on 1611 0.5 mW

  • In all three measurements, the phaselock was lost when the beat note drifted too close to DC. I'm trying again with the SLP-100 to give myself some more range.

  • really need to put these fibers in a box

While waiting for the measurements,

Update:

  2687   Thu Mar 18 18:21:36 2021 aaronDailyProgressLaser3CH setup

Set up the three corner hat in the same optical configuration as previously.

All three beat notes are on the same 1611-FC, then sent to a 4-way RF splitter. RF components we could use more of:

  • SMA 50 Ohm terminators
  • low pass filters above 50 MHz
  • BP filters wider than 6 MHz (or some HP to match the LP)
  • N to SMA adapter
  • 1->2 SMA terminated RF splitters

 

Splitter output sent to (fig pending, I have it hand-scrawled):

  1. 50 Ohms
  2. DC block -> SHP-100+ -> pomona box 135 MHz lowpass (47 pF cap to ground) -> mixer R (for marconi)
  3. SHP-150 -> Moku 2
  4. SLP-100 -> Moku 1

Struggled to get the Marconi locked, though I did manage a couple times... tomorrow.

 

  2686   Thu Mar 18 16:44:33 2021 shrutiDailyProgressPSOMAcavity alignment, RF oscillator power and frequency

Proceeded to align the cavity as outlined in elog 2685.

 

I also measured the output of OCXO using the Moku; all channels have 10dB attenuation and 50 Ohm impedance.

Attachment 1: 32.7 MHz; Channel 1 (red) is 'TO EOM', Channel 2 (blue) is 'TO LO'.

Attachment 2: 33.59 MHz; Channel 1 (red) is 'TO EOM', Channel 2 (blue) is 'TO LO'.

Attachment 3: Channel 1 (red) is 'TO LO' at 32.57 MHz, Channel 2 (blue) is 'TO LO' at 33.59 MHz.

The data files for all the spectra can be found here.

 

 

  2685   Thu Mar 18 10:08:40 2021 shrutiUpdatePSOMACavity design

Using the latest beam profiles in Attachments 3,4 and the design in Attachment 1, I calculated the mode-matching solutions found here. The one I decided to go with is in Attachment 2.

  2684   Wed Mar 17 12:55:42 2021 aaronDailyProgressstuff happensnot much

enter Wed Mar 17 12:55:44 2021

Grabbed the Marconi from cryo lab.

I left some of the current noise data on the moku ipad, so I transfer that over to spirou to plot the current noise.

Thought about filters for a while. Stuffed an RC low pass in a pomona box, but realized an LC would be better for the job.

exit Wed Mar 17 17:13:55 2021

  2683   Mon Mar 15 12:11:51 2021 shrutiNotesLaserPower vs. Temperature plots for Rio PLANEX

South Laser [Attachment 1]

The power dependence on temperature at different diode currents:

  • The plots are linear in sensor resistance on the x-axis
  • (Top figure) At 130 mA, while the sensor resistance decreased, an abrupt change was seen at 6.73 kOhms, and while increasing it was at 8.27 kOhms. At other regions it is locally reversible, i.e., going back and forth causes the same curve to be plotted.
  • (Top figure) This trend is repeatable, as the black lines show the results of the same change performed a second time.
  • (Middle figure) At around 120 mA, three such abrupt changes was seen in the range the data was taken.
  • (Bottom figure) Similar plot at 100 mA.

Data and notebook for plotting in Attachment 3

Conversion to Temperature

This notebook uses Rtyp values mentioned on the Rio Planex Datasheet found on the wiki documents page to obtain the plot and fit in Attachment 2.

That was used to make the temperature in deg C axes in Attachment 1.

  2682   Thu Mar 11 13:56:47 2021 aaronDailyProgressElectronicscurrent noise

enter Thu Mar 11 13:56:46 2021

  • On Busby low noise box AC coupled input (logged by to 1 MOhm moku channel 2)
    • measure voltage noise of a 20.5 Ohm resistor for 28 minutes
    • measure voltage noise across the same resistor in driven by ITC502 at 119.38 mA, same duration
    • python [path/to/moku]mokuDataLogger.py -c ch2 -i 10.0.5.253 -d 1680 -s 64e3 --useInternal --fileType bin --MOhmr2
    • started to make a box for sensing current noise... but perhaps this is not necessary. To measure 10 kHz/rtHz frequency noise on our 147 MHz/mA diode (-> uV/rtHz), even the metal film resistors (1% tol) with the steepest temperature coefficients on digikey (~100 ppm/K) could allow 10 mK/rtHz temperature fluctuations of the resistor. Still, the room is drafty, so I also measured the noise across the resistor driven by two AAA batteries (2.4 V total, for about the same current as before).
    • What is the voltage across the diode during operation, and can I check this at some of the butterfly pins? Does the current supply remain consistent operating at the same current but a different voltage? And, shouldn't a 'low noise current supply' be quieter than a battery (at least at the frequencies where it matters) anyway?
  • While waiting for the data, assembled the wire shelf on the West wall of PSOMA, between the vacuum cabinet and He cryostat.
    • Electronics on the 2nd shelf have rails for earthquake protection. We have sufficient rails for all 4 shelves, but I left them for another day.
    • Zip-tied the upper shelves to the piping on the wall. Seems sturdy.
    • Feet have plastic sliders to dampen the load on the floor. We also have label holders in need of labeling.
    • Photos are in the ligo.wbridge google drive and some attached.

exit Thu Mar 11 23:11:33 2021

 

  2681   Wed Mar 10 16:03:19 2021 aaronDailyProgressElectronicscurrent noise

enter Wed Mar 10 16:02:33 2021

I have "The Busby Low Noise Box" from the 40m and will be measuring current noise of our laser drivers... but today I mostly played around with liso. If anything useful, getting more accustomed to vectorizing python functions.

exit Wed Mar 10 21:07:52 2021

  2679   Tue Mar 9 14:03:12 2021 aaronComputingDAQoma model
  • installed cds-workstation on spirou (for some reason, was being held back by an incorrect version of python3-gpstime. Manually installing the correct version of that package let me apt install cds-workstation).
  • fixing bugs in x1oma model, following instructions on dcc.
    • Need at least 2 filter modules to build the model (it was just a dummy model with ADC and DAC).
    • 'site' designator is deprecated, I've replaced it with 'ifo=X1'
    • Need to use a different testpoint node (dcuid) from the existing models. 
    • from cymac1
      • rtcds build x1oma
        rtcds install x1oma
      • restart rtcds. The x1oma model wasn't added to the startup sequence or status screens.
        • Manually running rtcds start x1oma completes successfully, and /opt/rtcds/tst/x1/medm/x1oma contains the usual .adl files.
        • Now, after building and installing x1oma is in the rtcds startup sequence, but doesn't load successfully
        • Possibly just built the model with an incompatible version of matlab. Gaston is running 2015b, which works historically, and is also the latest Matlab I see mentioned in RCG release notes.
          • When building the x1oma model (with either Matlab 2020 or 2015 versions), I get a warning before the model eventually compiles successfuly
            • cat: '/opt/DIS/lib/modules/4.19.0-6-rtcds-amd64/*.symvers': No such file or directory
            • This warning is also there when I compile other models (like x1siq), even though they were built with Matlab 2015b.
          • Saving the model as a 2015b compatible model does not lead to an installed x1oma
        • I'm using some 'already in use' DAC channels in x1oma. In fact, the only 2 free DAC channels (of our 16 channel fast DAC) are 1 and 2. 0 is occupied by SIQ ESD; 3-7 occupied by Si cantilever Qs; and 8-16 by cryo cavs. I swapped the OMA to use DAC channels 1 and 2, but this x1oma still fails to load (and doesn't auto-populate status medm screens).
      • This has to be lower than the x1oma model. rtcds status shows several kernel modules are not loaded (dis_*).
        • On cymac1 I apt install advligorts-cymac, I'm told the package was not fully installed and it updates to 4.1.0-1 (was 4.0.1).
        • I confirmed that these modules are running on one of the 40m frontend machines.
        • I reinstalled cds-workstation and advligorts-cymac (no change)
        • rebuilt and reinstalled all models. No change, in particular status screens not updated with x1oma model.
          • Oddly, x1oma is no longer in the startup sequence or list of models...
Cymac isn't crashing, so I'll vacate the lab and continue from home
  2678   Tue Mar 9 13:57:38 2021 aaronDailyProgress update

enter Tue Mar 9 13:57:31 2021

  • picked up a fiber microscope and cleaner from Aidan's office
  • I'd like to record data with cymac1 to assist in the 'coherent 3 corner hat' measurement. To that end, I'm setting up the x1oma model (again).
    • Not much success

exit Tue Mar 9 16:42:41 2021

ELOG V3.1.3-